# Deep Learning; Physics-Informed Neural Networks; Partial Differential Equations; Schrödinger equation

### v1

@@ 3 How to evaluate training
In this section some ways to evaluate the PINN’s training in order to optimize their architecture and hyperparameters are shown. The focus will be on some typical pathological behaviour and some ways to understand what causes them and how to fix them. The behaviour of the loss actually tends to be a pretty reliable metric when training PINNs in an unsupervised way, and is actually more reliable then for standard neural networks. This is because a PINN follows a given physical law, and the loss quantifies how well its output follow said law. It does not have a set of training points it might memorize instead of learning the underlying law. As such, overfitting on some training set is not a risk and it does not make sense to build a training, validation and test sets from a dataset. In spite of this, due to the complexity of the loss function just looking at the full loss does not tell the whole story of what is happening during training. In the following subsections some ways to interpret the behavior of the network during training more deeply will be detailed. This should allow one make more informed decisions when tuning hyperparameters.
@@ 4.2 Particle in a ring
5 Conclusion
The application and development of deep learning for science and technology has huge potential, and PINNs are among the most promising tools to progress science by using machine learning. Nowadays, one of the most stimulating fields of research aims at the development of novel technologies able to actively leverage quantum resources. For such purposes, the characterization of the energy spectrum (and the corresponding eigenfunctions) of the system is essential and it passes through the solution of the Schrödinger’s equation. This tutorial provided the bases to build a PINN to solve the Schrödinger’s equation in an unsupervised way. We have discussed the infinite potential well and the particle in a ring, exemplifying systems that allow the reader to fully understand the essential steps to build such type of neural networks. The research in the use of PINNs in quantum technologies is still relatively novel, requiring more research to find the optimal way to employ such technique. Future directions include finding a better way to optimize the weights, improve on the computation of integrals, building models able to better handle degenerate states, and most importantly applying the method to increasingly more complex systems.
@@ 5 Conclusion
The Authors acknowledge the project CQES of the Italian Space Agency (ASI) for having partially supported this research (grant N. 2023-46-HH.0).

### v2

@@ 3 How to evaluate training
In this section, some ways to evaluate the PINN’s training to optimize their architecture and hyperparameters are shown. The focus will be on some typical pathological behaviors and some ways to understand what causes them and how to fix them. The behavior of the loss actually tends to be a pretty reliable metric when training PINNs in an unsupervised way, and is actually more reliable than for standard neural networks. This is because a PINN follows a given physical law, and the loss quantifies how well its output follows the given law. It does not have a set of training points, it might memorize instead of learning the underlying law.
As such, overfitting on some training sets is not a risk and it does not make sense to build the training, the validation and the test sets from a dataset. In spite of this, due to the complexity of the loss function just looking at the full loss does not tell the whole story of what is happening during training. In the following subsections, some ways to interpret the behavior of the network during training more deeply will be detailed. This should allow one to make more informed decisions when tuning hyperparameters.
@@ 4.2 Particle in a ring
5 Challenges, limitations and poential of PINNs
PINNs have the potential to be able to scale rather well with the complexity of the system. This is thanks to the extreme expressive power of Deep Neural Networks, which are known to be able to successfully approximate any continous function Cybenko (1989); Kim and Adalı (2003); Hornik (1991); Le Roux and Bengio (2008), given a sufficient amount of parameters. However, PINNs do have limitation and disadvantages compared to other computational methods. First of all, the multiple losses create an extremely complex loss landscape, which can be very frustrated Krishnapriyan et al. (2021). This makes training the PINN rather difficult, and makes it very likely for it to get stuck in a local minima if the hyperparameters have not been set correctly or if the initialization was especially unlucky. Overcoming this challenge is essential to make the method feasible for more complex systems. A way to do that would be to optimize the way weights are set, going beyond heuristics and trial-and-error Thanasutives et al. (2021); Xiang et al. (2022); Bu et al. (2020). Furthermore, for potentials with degenerate states, the PINN has no way to prefer any particular set of states, and is equally likely to output any linear combination of the degenerate states. This is generally not an issue if we are just looking for solutions of Schrödinger’s equation for the given potential, since the PINNs will output valid solutions. However, if a particular set of eigenfunctions is preferred, we need to add further physical constraint to the network to ensure the correct output. It is also important to note that for these simple systems the PINN’s performances are inferior to those of numerical solvers, both from the point of view of accuracy and speed, if we take into account the training time. For instance, numerical inversion of the Hamiltonian for the particle in a ring leads to errors for the eigenvalues in the order of $10^{-11}$ and can be performed in a few seconds. However, while there are still many challenges with PINNs, especially related to their training, this tool is potentially able to represent any quantum mechanical system given either the governing potential or a sufficient amount of data. Possibly even going beyond what current numerical solvers are able to do. Furthermore, a PINN trained on a more complex system could be utilized to quickly output the wave function $\psi(\vec{x})$ at any point $\vec{x}$, possibly faster then classical solvers.
6 Conclusions
The application and development of deep learning for science and technology has huge potential, and PINNs are among the most promising tools to progress science by using machine learning. Nowadays, one of the most stimulating fields of research aims at the development of novel technologies able to actively leverage quantum resources. For such purposes, the characterization of the energy spectrum (and the corresponding eigenfunctions) of the system is essential and it passes through the solution of the Schrödinger’s equation. This tutorial provided the bases to build a PINN to solve the Schrödinger’s equation in an unsupervised way. We have discussed the infinite potential well and the particle in a ring, exemplifying systems that allow the reader to fully understand the essential steps to build such type of neural networks. The research in the use of PINNs in quantum technologies is still relatively novel, requiring more research to find the optimal way to employ such technique. Future directions include finding a better way to optimize the weights, improve on the computation of integrals, building models able to better handle degenerate states, and most importantly applying the method to increasingly more complex systems, such as atomic nuclei, to assess the role played by coherence in quantum technologiesZhang et al. (2024), and to solve stochastic Schrödinger’s equation modelling decohering processes Chiarelli (2023).
@@ 6 Conclusions
The Authors acknowledge the project CQES of the Italian Space Agency (ASI) for having partially supported this research (grant N. 2023-46-HH.0). The authors also acknowledge support from the Qxtreme project funded via the Partenariato Esteso FAIR.